I hope that this is the right place to put this question. I feel like there isn't enough information given to me to solve it, but it's all I was given. I'm really hoping someone can at least help me get started.
In a college meal plan you pay a membership fee; then all your meals are at a fixed price per meal.
(a) If 34 meals cost $211.50 and 67 meals cost $368.25, find the membership fee and the price per meal.

Answer by richwmiller(9143)

You want to write a linear equation in the form y=mx+b.
The two points are (34,211.50) and (67,368.25)
first use the formula
(y-y)/(x-x)=m
(368.25-211.50)/(67-34)=m
y=mx+b
put the new m in the equation and either ordered pair in for x and y to find b
b will be the membership fee
m will be the fixed price per meal